OpenStudy (migitmack):

What is the simplified form of the expression 4√(v)+4√(v^6)-√(25v)-√(9v^6)? 6v^2 sqrt(v) v3 −sqrt(v) 6v^5 sqrt(v) sqrt(v-v^3)

OpenStudy (migitmack):

plz hellp me

OpenStudy (anonymous):

it's simple, remember that \[\sqrt{v^{2n}}=v^n, \sqrt{9} = 3, \sqrt{25}=5\] and that \[\sqrt{ab}=\sqrt{a}\sqrt{b}\], so you can simplify most of the terms